OpenStudy (anonymous):

In the neighborhood park there is a circular grassy picnic area, but half of the grass has died. The city wants to replant the grass, but needs to know the square footage of the damaged area. If the circular picnic area is 150 feet across, what is the square footage of the damaged area? Let π = 3.14. A)8831.25 sq. feet B)17662.5 sq. feet C)35325 sq. feet D)70650 sq. feet

OpenStudy (anonymous):

jljj\[\csc \iota \times \int\limits_{\sqrt[\tan^{-1} \left\{ ? \right\}]{?}}^{?}\] use this equation, its easy